# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v1

@@ 1.3 Broyden family of optimizers
Many quasi-Newton methods fall under the Broyden family of updates, which is characterized by the following general formula:
@@
where $\theta_{k}$ is a scalar parameter that could vary between iterations, and
$\mathbf{w}_{k}=\frac{\mathbf{y}_{k}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}}-\frac{\mathbf{B}_{k}\mathbf{s}_{k}}{\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k}}$
The BFGS and DFP methods are both special cases of the Broyden class. Specifically, we recover the BFGS method by setting $\theta_{k}=0$ and the DFP method by setting $\theta_{k}=1$. The search direction for quasi-Newton methods are defined as
$\mathbf{p}_{k}=-\mathbf{H}_{k}\nabla f_{k},$
where $\mathbf{H}_{k}$ denotes the inverse of $\mathbf{B}_{k}$. In practice, the different Quasi-Newton formulas to update the Hessian approximations are especially meant to work directly with $\mathbf{H}_{k}$, thus avoiding any matrix inversion during the calculations. Certainly, if we apply the inverse in the expression (10) at both sides, then with the aid of the well-known Sherman-Morrison formula in matrix theory, we get
@@
$\phi_{k}$ $=\frac{1-\theta_{k}}{1+\left(h_{k}b_{k}-1\right)\theta_{k}},$
@@
If this parameter depends explicitely on $\mathbf{B}_{k}$, then the expressions (10) and (11) should be employed in every iteration to update the inverse Hessian estimations. However, in the special case where this dependency on $\mathbf{B}_{k}$ appears via its product with $\mathbf{s}_{k}$, then we can avoid using (10), since
@@
Among the various methods included in the Broyden class, the BFGS method has demonstrated particular efficacy in solving small and medium-sized unconstrained optimization problems [57, 58]. However, numerical experience shows that its effectiveness drops considerably for certain ill-conditioned problems (see for example [59, 60]). To mitigate this drawback, the self-scaled BFGS (SSBFGS) method, firstly proposed by Oren and Luenberger [61], has proven to be one of the most promising optimization techniques. In SSBFGS, the Hessian estimate $\mathbf{B}_{k}$ is scaled by some scalar $\tau_{k}>0$, which can also vary with iterations, before updating it using equation (2.19). The motivation of this scaling was to reduce the condition number of the matrix $\mathbf{H_{k}}^{1/2}\nabla^{2}f_{k}\mathbf{H_{k}}^{1/2}$, which serves as a good indication of the rate of convergence for Quasi-Newton methods. Further theoretical and numerical analysis provided by Nocedal and Yuan in [62] yielded discouraging results for one of the best suggestions of Oren and Luenberger, defined as $\tau_{k}=1/b_{k}$. However, later on Al-Baali [63] showed encouraging results with the simple modification $\tau_{k}=\min\left\{1,1/b_{k}\right\}$. Al-Baali [64] extended the scaling technique to other members of the Broyden family of updates, demonstrating promising theoretical and computational results for various choices of $\theta_{k}\in\left[0,1\right]$, while previously imposing the constraint $\tau_{k}\leq 1$. Additional families of scaling methods for BFGS and other Broyden family updates—presented in works such as [65, 66, 67, 68, 69]—further support the effectiveness of scaling techniques for certain optimization problems. The self-scaled updates for the direct and the inverse hessian estimates can be easily extended for the general Broyden family, which will be called SSBroyden updates in the following. These are given by
$\mathbf{B}_{k+1}$ $=\tau_{k}\left[\mathbf{B}_{k}-\frac{\mathbf{B}_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{\top}\mathbf{B}_{k}}{\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k}}+\theta_{k}(\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k})\mathbf{w}_{k}\mathbf{w}_{k}{\top}\right]+\frac{\mathbf{y}_{k}\mathbf{y}_{k}^{\top}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}},$ (13)
$\mathbf{H}_{k+1}$ $=\frac{1}{\tau_{k}}\left[\mathbf{H}_{k}-\frac{\mathbf{H}_{k}\mathbf{y}_{k}\mathbf{y}_{k}^{\top}\mathbf{H}_{k}}{\mathbf{y}_{k}^{\top}\mathbf{H}_{k}\mathbf{y}_{k}}+\phi_{k}(\mathbf{y}_{k}^{\top}\mathbf{H}_{k}\mathbf{y}_{k})\mathbf{v}_{k}\mathbf{v}_{k}^{\top}\right]+\frac{\mathbf{s}_{k}\mathbf{s}_{k}{\top}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}}.$ (14)
@@
In PINNs, specifically using Multi-Layer Perceptrons (MLPs), recent studies have explored different families of SSBroyden updates and modifications to the loss function. Urbán et al. [50] first analyzed these approaches in detail, demonstrating that for various types of PDEs, these modifications significantly improve both convergence and accuracy compared to BFGS. In this work, we extend these findings to more complex problems beyond those presented in [50], employing not only SSBroyden but also other techniques proposed in the PINN literature. Additionally, we investigate whether these BFGS modifications enhance the performance of the recently introduced Kolmogorov-Arnold Networks (KANs) [70]. This study will further enable fair comparisons between KANs and MLPs using more advanced optimization algorithms. In the following section, we present a comprehensive study comparing the performance of different optimization methods with various line-search and trust-region strategies for solving PDEs. For each example, we analyze and discuss the performance of both PINNs and PIKANs.
@@ 2.1.1 BFGS and SSBroyden optimization with Strong Wolfe line-search
A comparative study of the performance of L-BFGS, BFGS and SSBroyden optimization with Strong Wolfe line-search for solving the Burgers’ Equation has been conducted using seven different case studies. In all PINNs cases, the hidden layers utilize the hyperbolic tangent ($\tanh$) activation function. Model training is performed in two stages: first, the Adam optimizer is employed with a learning rate of $10^{-3}$ for 1000 iterations. Moreover, the adaptive sampling strategy (RAD) algorithm defined in [71] is employed to dynamically resample points in regions with high errors.
@@
1 Adam [1000] + BFGS with Wolfe [10000] $1.04\times 10^{-4}$ 488 1,341
1 Adam [1000] + SSBroyden with Wolfe [10000] $3.60\times 10^{-5}$ 437 1,341
@@
Table 2: Double-Precision PINNs for the Burgers equation: Case 2: A PINN with four layers, each containing 20 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 50,000 iterations with BFGS, SSBroyden, and L-BFGS optimizers with Strong Wolfe line-search. Case 3: The same network structure as Case 2 is used, but the number of iterations with BFGS and SSBroyden is reduced to 30,000. Case 4: The same network structure as in Case 2 and Case 3 is used; however, instead of starting with the Adam optimizer, the training begins directly with second-order optimizers, namely BFGS and SSBroyden. Case 5: A deeper PINN with eight layers (20 neurons per layer) is trained for 1000 iterations using Adam, followed by 30,000 iterations with BFGS and SSBroyden.
@@
Figures 4 illustrate the loss function and relative error over iterations for Case 1 in single precision. The results are obtained using a PINN with four layers, each comprising 20 neurons, trained with the Adam optimizer followed by 10,000 iterations of the specified optimizer. Table 1 provides a summary of the corresponding relative error and training time. Notably, SSBroyden achieves a relative error of $10^{-5}$, while BFGS converges to approximately $10^{-4}$. It is observed that the loss function and relative error stabilize after roughly 4,000 iterations, indicating no significant further improvement.
The performance of PINNs with a similar network architecture and double precision is presented in Table 2 for Case 2 and Case 3. The same network architecture is employed, consisting of four hidden layers, each with 20 neurons. After the initial Adam optimization stage, the training proceeds with BFGS and SSBroyden optimizers for two different iteration counts: Case 2 utilizes 50,000 iterations, while Case 3 uses 30,000 iterations. In Case 4, the same network structure as in Case 2 and Case 3 is used; however, training begins directly with second-order optimizers, BFGS and SSBroyden, instead of the Adam optimizer. The results, as shown in the table, demonstrate that increasing the iteration count to 50,000 significantly raises the training time. While this reduces SSBroyden’s error to $10^{-8}$, it does not result in any noticeable improvement in the error achieved by BFGS. In Case 5, a deeper network with eight hidden layers is used, trained for 30,000 iterations. This configuration further improves the error, though at the cost of a substantial increase in training time. Figure 5 illustrates the loss function and relative error over iterations for Case 2 using three optimizers, BFGS, SSBroyden, and L-BFGS. As summarized in Table 2, the relative errors achieved are $10^{-8}$ for SSBroyden, $10^{-5}$ for BFGS, and $10^{-3}$ for L-BFGS.
@@ 3 Summary
In this study, we conducted a comprehensive comparison of multiple second-order optimizers for PINNs and PIKANs, focusing on BFGS, SSBroyden, and L-BFGS with Wolfe line-search conditions, as well as BFGS with Backtracking line search and trust-region methods. Utilizing the optax and optimistix library in JAX, we systematically evaluated these optimization techniques across a range of PDEs. We note that we obtained state-of-the-art results by only employing the bets optimizers without fine-tuning our loss functions with self-adaptive or attention-based weights or any other enhancemts. We simply pursued a straightforward application of good optimization solvers in order to demonstrate that the big bottleneck for PINNs or PIKANs is the optimization error.
@@
Our initial investigation focused on the Burgers equation, where we assessed the impact of combining first-order optimizers such as Adam with second-order optimizers like BFGS and SSBroyden. We then extended our analysis to more challenging PDEs, including the Allen-Cahn equation, the Kuramoto-Sivashinsky equation, and the Ginzburg-Landau equation, providing a detailed evaluation of efficiency and accuracy. The results demonstrate that advanced quasi-Newton methods, particularly SSBroyden and BFGS with Wolfe line-search, significantly improve the convergence rate and accuracy of PINNs, especially for complex and stiff PDEs. Across all cases, SSBroyden consistently outperformed BFGS, achieving faster convergence and exhibiting greater robustness in handling complex optimization landscapes. These findings highlight the effectiveness of quasi-Newton methods in accelerating the training of PINNs and enhancing numerical stability. Furthermore, we extended our analysis to PIKANs, replacing traditional multilayer perceptron architectures with KANs utilizing Chebyshev polynomials. In this setting, SSBroyden continued to demonstrate superior optimization performance, reinforcing its robustness across diverse network architectures. Additionally, our study emphasized the importance of implementing PINNs and PIKANs with double-precision arithmetic, which improves numerical stability and enhances optimization efficiency across all tested scenarios. This study demonstrates that SSBroyden with Wolfe line-search is an effective and reliable optimizer for training PINNs and PIKANs. Its capacity to handle complex PDEs and its strong convergence properties make it well-suited for advanced applications in deep learning-based scientific computing. Future research could focus on incorporating domain-specific adaptations to further improve the accuracy, efficiency, and generalization of PINNs and PIKANs, particularly in high-dimensional and multi-scale problem settings.

### v2

@@ 1.3 Broyden Family of Optimizers
Many quasi-Newton methods fall under the Broyden family of updates, characterized by the following general formula:
@@
where $\theta_{k}$ is a scalar parameter that may vary at each iteration, and
$\mathbf{w}_{k}=\frac{\mathbf{y}_{k}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}}-\frac{\mathbf{B}_{k}\mathbf{s}_{k}}{\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k}}.$
The BFGS and DFP methods are special cases of the Broyden class. Specifically, setting $\theta_{k}=0$ recovers the BFGS update, while $\theta_{k}=1$ yields the DFP update. In practice, quasi-Newton methods typically work directly with $\mathbf{H}_{k}$, avoiding explicit matrix inversion. By applying the inverse to both sides of equation (10), and using the Sherman–Morrison formula, we obtain the following update for $\mathbf{H}_{k}$:
@@
$\phi_{k}$ $=\frac{1-\theta_{k}}{1+(h_{k}b_{k}-1)\theta_{k}},$
@@
Additional useful expressions include:
$a_{k}$ $=b_{k}h_{k}-1,$
$c_{k}$ $=\left(\frac{a_{k}}{1+a_{k}}\right)^{1/2},$
$\rho_{k}^{-}$ $=\min\left(1,h_{k}(1-c_{k})\right),$
$\theta_{k}^{-}$ $=\frac{\rho_{k}^{-}-1}{a_{k}},$
$\theta_{k}^{+}$ $=\frac{1}{\rho_{k}^{-}},$
$\theta_{k}$ $=\max\left(\theta_{k}^{-},\min\left(\theta_{k}^{+},\frac{1-b_{k}}{b_{k}}\right)\right),$
$\rho_{k}^{+}$ $=\min\left(1,\frac{1}{b_{k}}\right),$
$\sigma_{k}$ $=1+\theta_{k}a_{k},$
$\sigma_{k}^{(1-N)}$ $=|\sigma_{k}|^{\frac{1}{1-N}},$
$\tau_{k}$ =$=\begin{cases}\min\left(\rho_{k}^{+}\sigma_{k}^{(1-N)},\sigma_{k}\right),&\text{if }\theta_{k}\leq 0,\\[6.0pt] \rho_{k}^{+}\min\left(\sigma_{k}^{(1-N)},\frac{1}{\theta_{k}}\right),&\text{otherwise}.\end{cases}$
If $\theta_{k}$ depends explicitly on $\mathbf{B}_{k}$, then both equations (10) and (11) must be used at each iteration to update the inverse Hessian estimates. However, when $\mathbf{B}_{k}$ appears only through the product $\mathbf{B}_{k}\mathbf{s}_{k}$, the update can avoid direct use of (10), since:
@@
Among the methods in the Broyden class, the BFGS method is particularly effective for small- and medium-scale unconstrained optimization problems [62, 63]. However, its performance may deteriorate for ill-conditioned problems [64, 65]. To address this limitation, the self-scaled BFGS (SSBFGS) method was proposed by Oren and Luenberger [66]. In SSBFGS, the Hessian approximation $\mathbf{B}_{k}$ is scaled by a positive scalar $\tau_{k}$ prior to the BFGS update, i.e.,
$\mathbf{B}_{k+1}=\tau_{k}\left[\mathbf{B}_{k}-\frac{\mathbf{B}_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{\top}\mathbf{B}_{k}}{\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k}}\right]+\frac{\mathbf{y}_{k}\mathbf{y}_{k}^{\top}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}}.$ (13)
The motivation behind scaling is to reduce the condition number of $\mathbf{H}_{k}^{1/2}\nabla^{2}f_{k}\mathbf{H}_{k}^{1/2}$, which reflects the convergence rate. While initial results by Nocedal and Yuan [67] were discouraging for the scaling $\tau_{k}=1/b_{k}$, later work by Al-Baali [68] showed promising results using the modified scaling $\tau_{k}=\min\{1,1/b_{k}\}$. Al-Baali [69] extended the idea of scaling to other Broyden updates, proving favorable theoretical and numerical results for various $\theta_{k}\in[0,1]$, while previously assuming $\tau_{k}\leq 1$. Additional scaling strategies have been developed for BFGS and other Broyden methods in works such as [70, 71, 72, 73, 74]. The self-scaled versions of the direct and inverse updates for general Broyden family methods—termed SSBroyden methods—are:
$\mathbf{B}_{k+1}$ $=\tau_{k}\left[\mathbf{B}_{k}-\frac{\mathbf{B}_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{\top}\mathbf{B}_{k}}{\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k}}+\theta_{k}(\mathbf{s}_{k}^{\top}\mathbf{B}_{k}\mathbf{s}_{k})\mathbf{w}_{k}\mathbf{w}_{k}^{\top}\right]+\frac{\mathbf{y}_{k}\mathbf{y}_{k}^{\top}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}},$ (14)
$\mathbf{H}_{k+1}$ $=\frac{1}{\tau_{k}}\left[\mathbf{H}_{k}-\frac{\mathbf{H}_{k}\mathbf{y}_{k}\mathbf{y}_{k}^{\top}\mathbf{H}_{k}}{\mathbf{y}_{k}^{\top}\mathbf{H}_{k}\mathbf{y}_{k}}+\phi_{k}(\mathbf{y}_{k}^{\top}\mathbf{H}_{k}\mathbf{y}_{k})\mathbf{v}_{k}\mathbf{v}_{k}^{\top}\right]+\frac{\mathbf{s}_{k}\mathbf{s}_{k}^{\top}}{\mathbf{y}_{k}^{\top}\mathbf{s}_{k}}.$ (15)
@@
In the context of Physics-Informed Neural Networks (PINNs) based on Multi-Layer Perceptrons (MLPs), recent research has begun to explore various families of SSBroyden updates and modifications to the loss formulation to address known challenges such as gradient conflicts and ill-conditioning. These second-order strategies significantly enhance both convergence speed and prediction accuracy compared to classical BFGS. In this work, we extend these findings to more complex problems beyond those presented in [53], employing not only SSBroyden but also other techniques proposed in the PINN literature. Additionally, we investigate whether these BFGS modifications enhance the performance of the recently introduced Kolmogorov-Arnold Networks (KANs) [3]. This study will further enable fair comparisons between KANs and MLPs using more advanced optimization algorithms. In the following section, we present a comprehensive study comparing the performance of different optimization methods with various line-search and trust-region strategies for solving PDEs. For each example, we analyze and discuss the performance of both PINNs and PIKANs.
@@ 2.1.1 BFGS and SSBroyden optimization with Strong Wolfe line-search
A comparative study of the performance of L-BFGS, BFGS, SSBFGS and SSBroyden optimization with Strong Wolfe line-search for solving the Burgers’ Equation has been conducted using seven different case studies. In all PINNs cases, the hidden layers utilize the hyperbolic tangent ($\tanh$) activation function. Model training is performed in two stages: first, the Adam optimizer is employed with a learning rate of $10^{-3}$ for 1000 iterations. Moreover, the adaptive sampling strategy (RAD) algorithm defined in [75] is employed to dynamically resample points in regions with high errors.
@@
1 Adam [1000] + BFGS with Wolfe [10000] $1.04\times 10^{-4}$ 263 1,341
1 Adam [1000] + SSBroyden with Wolfe [10000] $3.60\times 10^{-5}$ 209 1,341
1 Adam [1000] + SSBFGS with Wolfe [10000] $4.04\times 10^{-5}$ 179 1,341
@@
5 Adam [1000] + SSBFGS with Wolfe [30000] $4.19\times 10^{-8}$ 2878 3,021
Table 2: Double-Precision PINNs for the Burgers equation: Case 2: A PINN with four layers, each containing 20 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 50,000 iterations with BFGS, SSBroyden, SSBFGS, and L-BFGS optimizers with Strong Wolfe line-search. Case 3: The same network structure as Case 2 is used, but the number of iterations with BFGS, and SSBroyden is reduced to 30,000. Case 4: The same network structure as in Case 2 and Case 3 is used; however, instead of starting with the Adam optimizer, the training begins directly with second-order optimizers, namely BFGS, SSBFGS, and SSBroyden. Case 5: A deeper PINN with eight layers (20 neurons per layer) is trained for 1000 iterations using Adam, followed by 30,000 iterations with BFGS and SSBroyden.
@@
6 Adam [1000] + SSBFGS with Wolfe [50000] (degree 3) $1.77\times 10^{-8}$ 12796 5,040
@@
7 Adam [1000] + SSBFGS with Wolfe [50000] (degree 3) $4.21\times 10^{-6}$ 810 920
@@
Figures 4 illustrate the loss function and relative error over iterations for Case 1 in single precision. The results are obtained using a PINN with four layers, each comprising 20 neurons, trained with the Adam optimizer followed by 10,000 iterations of the specified optimizer. Table 1 provides a summary of the corresponding relative error and training time. Notably, SSBFGS and SSBroyden achieve a relative error of $10^{-5}$, while BFGS converges to approximately $10^{-4}$. It is observed that the loss function and relative error stabilize after roughly 4,000 iterations, indicating no significant further improvement.
The performance of PINNs with a similar network architecture and double precision is presented in Table 2 for Case 2 and Case 3. The same network architecture is employed, consisting of four hidden layers, each with 20 neurons. After the initial Adam optimization stage, the training proceeds with BFGS, SSBFGS, and SSBroyden optimizers for two different iteration counts: Case 2 utilizes 50,000 iterations, while Case 3 uses 30,000 iterations. In Case 4, the same network structure as in Case 2 and Case 3 is used; however, training begins directly with second-order optimizers, BFGS, SSBFGS, and SSBroyden, instead of the Adam optimizer. The results, as shown in the table, demonstrate that increasing the iteration count to 50,000 significantly raises the training time. While this reduces SSBFGS and SSBroyden’s error to $10^{-8}$, it does not result in any noticeable improvement in the error achieved by BFGS. In Case 5, a deeper network with eight hidden layers is used, trained for 30,000 iterations. This configuration further improves the error, though at the cost of a substantial increase in training time. Figure 5 illustrates the loss function and relative error over iterations for Case 2 using four optimizers, BFGS, SSBFGS, SSBroyden, and L-BFGS. As summarized in Table 2, the relative errors achieved are $10^{-8}$ for SSBroyden, $10^{-5}$ for BFGS, and $10^{-3}$ for L-BFGS.
@@ 3 Summary
In this study, we conducted a comprehensive comparison of multiple second-order optimizers for PINNs and PIKANs, focusing on BFGS, SSBFGS, SSBroyden, and L-BFGS with Wolfe line-search conditions, as well as BFGS with Backtracking line search and trust-region methods. Utilizing the optax and optimistix library in JAX, we systematically evaluated these optimization techniques across a range of PDEs. We note that we obtained state-of-the-art results by only employing the best optimizers without fine-tuning our loss functions with self-adaptive or attention-based weights or any other enhancemts. We simply pursued a straightforward application of good optimization solvers in order to demonstrate that the big bottleneck for PINNs or PIKANs is the optimization error.
@@
Our initial investigation focused on the Burgers equation, where we evaluated the impact of combining first-order optimizers such as Adam with second-order optimizers such as BFGS and SSBroyden. We then extended our analysis to more challenging PDEs, including the Allen-Cahn equation, the Kuramoto-Sivashinsky equation, and the Ginzburg-Landau equation, providing a detailed evaluation of efficiency and accuracy. The results demonstrate that advanced quasi-Newton methods, particularly SSBroyden, SSBFGS and BFGS with Wolfe line-search, significantly improve the convergence rate and accuracy of PINNs, especially for complex and stiff PDEs. Across all cases, SSBroyden consistently outperformed SSBFGS and BFGS, achieving faster convergence and exhibiting greater robustness in handling complex optimization landscapes. These findings highlight the effectiveness of quasi-Newton methods in accelerating the training of PINNs and enhancing numerical stability. Furthermore, we extended our analysis to PIKANs, replacing traditional multilayer perceptron architectures with KANs utilizing Chebyshev polynomials. In this setting, SSBroyden continued to demonstrate superior optimization performance, reinforcing its robustness across diverse network architectures. Additionally, our study emphasized the importance of implementing PINNs and PIKANs with double-precision arithmetic, which improves numerical stability and enhances optimization efficiency across all tested scenarios.
In Appendix B, we provide a detailed comparison of the number of iterations required to minimize the Rosenbrock function using various optimizers. This analysis highlights how increasing the dimensionality affects the convergence behavior of second-order methods such as BFGS and SSBroyden. We observed that these methods remain efficient across dimensions, while first-order optimizers like ADAM require significantly more iterations and fail to achieve comparable accuracy. The errors reported in this study show that using SSBroyden with the Wolfe line search leads to significantly lower relative $\mathcal{L}_{2}$ errors in all benchmark problems tested compared to the state-of-the-art SOAP method [56]. For instance, in the Burgers equation, our approach reduces the error from $8.06\times 10^{-6}$ to $4.19\times 10^{-8}$. In the Kuramoto–Sivashinsky equation, we achieve an error of $2.65\times 10^{-5}$ over the full time interval $t\in[0,1]$, whereas SOAP reports $3.86\times 10^{-2}$ over a shorter interval $t\in[0,0.8]$. Similar improvements are observed for the Allen–Cahn and Ginzburg–Landau equations (see Table 16).
Our study demonstrates that SSBFGS and SSBroyden with Wolfe line-search are effective and reliable optimizers for training PINNs and PIKANs. Their capacity to handle complex PDEs and their strong convergence properties make them well-suited for advanced applications in scientific machine learning applications. Future research should focus on incorporating domain-specific adaptations to further improve the accuracy, efficiency, and generalization of PINNs and PIKANs, particularly in high-dimensional and multi-scale problem settings.
Benchmark $\mathcal{L}_{2}$ Error (SOAP [56]) $\mathcal{L}_{2}$ Error (SSbroyden with Wolfe)
Burgers Equation $8.06\times 10^{-6}$ $4.19\times 10^{-8}$
Allen–Cahn Equation $3.48\times 10^{-6}$ $9.43\times 10^{-7}$
Kuramoto–Sivashinsky Equation $3.86\times 10^{-2}$ on $t\in[0,0.8]$ $2.65\times 10^{-5}$ on $t\in[0,1]$
Ginzburg–Landau Equation $4.78\times 10^{-3}$ $2.33\times 10^{-3}$
Table 16: Benchmark comparison of relative $\mathcal{L}_{2}$ errors. Results from SSBFGS with Wolfe line-search are compared against SOAP [56]. SSBFGS demonstrates improved accuracy across all PDE benchmarks.